# Coordinated Disclosure for AI: Beyond Security Vulnerabilities

### v2

@@ 3.3 Participatory Approaches for Issue Discovery
In response to growing public scrutiny around AI ethics and recognition of the severe limitations of ad hoc reporting for machine learning systems, some organizations have recently started experimenting with participatory approaches to systematically discover model issues.
@@
Bias bounty programs have emerged as one mechanism enabling security researchers and external auditors to uncover harmful biases and other flaws in ML models. These programs incentivize crowd-sourced identification of algorithmic problems by providing monetary rewards and public recognition for impactful findings. Some prominent early efforts in this space include Twitter’s algorithmic cropping bounty after concerns over racial bias (Yee, Tantipongpipat, and Mishra 2021) and QueerInAI’s queer bias bounty (Dennler et al. 2023) targeting intersectional issues, but AI bug bounties are slowly becoming commonplace in big companies as well (Google 2023; Microsoft 2023).
@@
Additionally, community-driven audits (Cattell, Chowdhury, and Carson 2023; The New York Times 2023), petitions (CAIR 2023), and pressure campaigns (Bloomberg 2023) have proven effective for advocacy groups to highlight urgent ML issues and hold organizations accountable. Truly participatory approaches, however, require extensive investments towards inclusion to lower barriers for participation by directly affected groups (Dennler et al. 2023). They also necessitate extensive transparency into model development and deployment choices to foster informed, constructive feedback. In addition, these approaches presently tend to have the authorization of the ML model owners, which the ML model owners could withhold, rescind, or condition to silence potentially damaging disclosures under the threat of a lawsuit (Friedler et al. 2023). Depending on ML model owners for authorization to research the ML models can ultimately undermine researchers’ independence.
@@
There are some nascent efforts towards standardized documentation of AI harms. Taxonomies such as MITRE’s ATLAS [^2] https://atlas.mitre.org/ and databases such as the AI Vulnerability Database (AVID) [^3] https://avidml.org/ managed by AI Risk and Vulnerability Alliance (ARVA), and the AI Incident Database (AIID) [^4] https://incidentdatabase.ai/ managed by the Responsible AI Collaborative are the most well known efforts in the field. AIID, well-established, collects news stories about AI incidents causing harm, while AVID assesses several open-source models and posts their findings on their database. Although useful, these initiatives lack coordinated disclosure features and mechanisms for vendor rebuttal. There are bounty programs like Google’s and Microsoft’s that demonstrate an embrace of disclosure at large model vendors (Google 2023; Microsoft 2023).
@@
AI red teaming has emerged as one of the most promising recent developments in surfacing ethical issues in machine learning before model launches (OpenAI 2023; Ganguli et al. 2022). It refers to an organization commissioning a diverse group of outside domain experts to rigorously test and attempt to attack their ML systems, simulating the behaviors of real-world malicious actors (Carey and Jin 2019; The White House 2023b).
The goals of AI red teaming exercises are to pressure test models, infrastructures, and deployment environments in adversarial ways, surface unknown issues or vulnerabilities that may have been overlooked during in-house testing, provide external validation of internal safety claims, and gain new perspectives for strengthening systems before public release (Selbst 2021; Moss et al. 2021; Carlini et al. 2023; Zou et al. 2023). Effective red teams require careful screening for a diverse mix of professional backgrounds, ideologies, regional representations, and skill sets.
While promising, these community-centered mechanisms remain the exception rather than the norm. Much work lies ahead in formalizing scalable, sustainable, and equitable participatory practices for ML issue discovery and disclosures.
@@ 4.1 Complete Model/System Card
The vendor needs to supply a model card for the complete system that includes intent and scope. These are related; intent covers what the system is meant to do, and scope limits the issues people can look for. This is already called for in the original paper (Mitchell et al. 2019) and implemented in many cards, including the Llama2 card (Meta 2023). Intent includes a mission statement for the system and the efficacy measures the model is tied to. These will evolve over time as the model is better understood in its environment. Scope is focused on known issues that the model has that are either in the process of being addressed or cannot be resolved in a reasonable timeframe. These irresolvable flaws come in two flavors: fundamental flaws in the system that can never be resolved and known flaws that are being worked on but might take a few years.
Essentially, this breaks up the idea of “scope” from the security community into two terms. This is needed as the idea of fallibility is built into our use of the term “intent,” but “in-scope” for security professionals means any vulnerabilities found in that scope. In security, the intent is often understood by the product description and community understanding. SQL injections are not a flaw in databases unless that database has a permissions scheme that the query was able to violate. As identical models, in terms of the input and output formats, can have wildly different intents, we believe that an intent statement will clear the air and reduce the amount of scope details needed. The Microsoft bounty program makes this distinction clear for LLMs, excluding prompt injections unless it can get privileged information from other users.
This model card is a complete system card. This system can contain known vulnerable models and additional models that are meant to mitigate those vulnerabilities. For example, using a diffusion model with an image net classifier might be secure against adversarial examples, and the model intent and scope can clearly state this as a system property where it is not a property of the individual model. This would make adversarial examples “in-scope” and reportable. Another example, for an LLM, you can deploy a hate speech model with the LLM that blocks hate speech like with NVIDIA Guardrails (Rebedea et al. 2023). The LLM model card should include “this potentially produces hate speech,” but the system can say “this does not produce hate speech.” Another example is Dalle2 inserting demographic descriptors into prompts to increase the diversity of the produced output (OpenAI 2022).
@@ 4.3 Automatic Verification
Errors in the model are going to be a large source of reports to any database of model issues. Reporting these to the vendor through a defined reporting system, a telemetry system for example, is preferable for the vendor for several reasons, but most importantly it enables easy integration into the model’s training and testing dataset. It also verifies that the model actually has the behavior and it’s reproducible, which may change day-to-day with model updates and continuous training. An outside report requires labor from the vendor to verify the behavior and add the input/output pair into the dataset.
Almost all useful reports listed in the background section include samples that demonstrate the behavior. Any system that aids in Coordinated Flaw Disclosure with the vendor should include support for automatic external verification and automatic integration in the vendors internal testing dataset. This significantly lowers the costs to the vendor to potentially remediate the issue, and provides a means of checking that the issue stays closed. If the issue is accepted the system should retain the data used to support the report and periodically re-verify the issue. If it is resolved by the vendor in one model, a subsequent model may regress and re-open the issue.
@@ 6 Conclusion
In this paper, we have proposed a Coordinated Flaw Disclosure (CFD) framework tailored specifically for machine learning and artificial intelligence systems. Our framework addresses the nuanced challenges posed by algorithmic flaws, emphasizing the importance of transparent intent and scope definitions through an extension of Model Cards for complete systems. By incorporating a trusted independent adjudication panel, we aim to facilitate fair dispute resolution, effectively distinguishing valid violations from issues rooted in sampling biases. The introduction of an automatic verification mechanism streamlines reporting processes, enhancing efficiency and ensuring issue reproducibility. With a structured issue reporting governance model encompassing triage, review, and adjudication, the CFD framework seeks to strike a balance between submitter and vendor interests, fostering trust and transparency. We believe that widespread adoption of this framework can contribute to bolstering public trust in AI systems by promoting clear processes and collaborative engagement between organizations and the broader community. As the field of machine learning evolves, the CFD framework provides a structured and adaptable approach to navigate the ever-changing landscape of algorithmic flaws, benefiting both vendors and the public alike.
@@ 6.1 Limitations
Model Cards need Standards
A notable limitation in implementing a CFD process is the lack of standardized model cards. Existing system card formats exhibit significant variation, and many models utilized by vendors lack public model cards. Addressing this discrepancy is crucial for widespread CFD adoption. We echo calls from the NIST AI Risk Management Framework for the establishment of standards for model cards (Singh 2021), providing a necessary foundation for developers, especially those with commercial ambitions.
Challenges of Dynamic ML Datasets
Another challenge stems from the complex nature of models built on diverse datasets, each potentially requiring a data card (Pushkarna, Zaldivar, and Kjartansson 2022). Drawing a parallel from the software domain, the concept of a Software Bill of Materials (SBOM) (CISA 2021) has been instrumental in regulating version control for software components (Corman 2022), essential for avoiding CVEs. However, the absence of equivalent tools or practices for tracking Dataset cards poses a significant hurdle. Unlike the well-established open-source tools for determining software dependencies, tracking the datasets used in model training, which evolve over time, presents a more intricate problem. Recent incidents, such as the retraction of the LAION-5B dataset (Schuhmann et al. 2022) due to the presence of Child Sexual Abuse Material (CSAM) (LAION.ai 2023; Thiel 2023), underscore the pressing need for robust mechanisms to manage dataset-related challenges in the machine learning landscape.
Common Weaknesses Enumeration: Contextual Challenges
Introducing a Common Weaknesses Enumeration (CWE) system for machine learning introduces novel challenges. Even for vulnerabilities, the weakness enumeration is highly context-specific. For example, a system might be theoretically vulnerable to the ATLAS (MITRE 2024b) weakness “Prompt Injection,” but due to the application’s design, this may not be a legitimate concern. In traditional software, a memory weakness like a buffer overflow in a C program is consistently a concern, often leading to vulnerabilities allowing remote code execution. While context dependency is not novel for AI systems, the introduction of identical ML models for very different purposes adds a layer of complexity. Unlike most traditional software contexts, the documentation around the model becomes crucial, introducing a dependency not typically encountered in conventional software scenarios.

### v3

@@ 3.3 Participatory Approaches for Issue Discovery
In response to growing scrutiny around AI ethics and the limitations of ad hoc reporting, organizations are experimenting with participatory approaches to systematically discover model issues.
@@
Bias bounty programs incentivize crowd-sourced identification of algorithmic problems by offering rewards and recognition for impactful findings. Early efforts include Twitter’s algorithmic cropping bounty (Yee, Tantipongpipat, and Mishra 2021) and QueerInAI’s queer bias bounty (Dennler et al. 2023). AI bug bounties are becoming common in large companies (Google 2023; Microsoft 2023).
@@
Community-driven audits, petitions, and pressure campaigns are effective for advocacy groups to highlight urgent ML issues and hold organizations accountable (Cattell, Chowdhury, and Carson 2023; The New York Times 2023). Truly participatory approaches require investments in inclusion to lower barriers for directly affected groups (Dennler et al. 2023). Transparency in model development and deployment fosters constructive feedback. However, relying on ML model owners for authorization can undermine researchers’ independence (Friedler et al. 2023).
@@
Standardized documentation of AI harms is nascent. Efforts like MITRE’s ATLAS [^2] https://atlas.mitre.org/, AI Vulnerability Database (AVID) [^3] https://avidml.org/, and AI Incident Database (AIID) [^4] https://incidentdatabase.ai/ are well-known. AIID collects news stories about AI incidents, while AVID assesses open-source models. These initiatives lack coordinated disclosure features and vendor rebuttal mechanisms. Bounty programs like Google’s and Microsoft’s show an embrace of disclosure at large model vendors (Google 2023; Microsoft 2023).
@@
AI red teaming involves organizations commissioning external experts to rigorously test and attempt to attack their ML systems, simulating real-world malicious actors (OpenAI 2023; Ganguli et al. 2022). Goals include pressure-testing models, surfacing overlooked issues, validating internal safety claims, and gaining new perspectives to strengthen systems before public release (Selbst 2021; Moss et al. 2021; Carlini et al. 2023; Zou et al. 2023). Effective red teams require diverse professional backgrounds, ideologies, regional representations, and skill sets.
While promising, these community-centered mechanisms are not yet the norm. Much work remains in formalizing scalable, sustainable, and equitable participatory practices for discovering and disclosing ML issues.
@@ 4.1 Complete Model/System Card
The vendor is required to provide a comprehensive model card for the entire system, detailing both the intent and scope. The intent encompasses the system’s purpose and objectives, while the scope delineates the limitations and known issues associated with the system. This requirement aligns with the guidance established in the seminal work by Mitchell et al. (Mitchell et al. 2019) and is exemplified in various model cards, such as the Llama2 card (Meta 2023).
The intent section should include a clear mission statement for the system and outline the efficacy measures relevant to the model. These elements are expected to evolve as the model’s performance and its contextual understanding improve. The scope section should address both current issues being addressed and those that are unresolved within a reasonable timeframe. These unresolved issues may be categorized into fundamental flaws that cannot be rectified and known flaws that are under ongoing development but may take several years to resolve.
This approach effectively distinguishes between two aspects of “scope” commonly used in the security domain. While “intent” encompasses the inherent fallibility of the system, “scope” in security contexts often refers to vulnerabilities within the specified parameters. The intent is generally understood through product descriptions and community feedback. Given that models with identical input and output formats may serve different purposes, a clearly defined intent statement helps clarify the system’s objectives and minimizes the need for extensive scope details. For instance, the Microsoft bounty program illustrates this distinction by excluding prompt injections from scope unless they access privileged information from other users.
A complete system card encompasses all components of the system, including both vulnerable models and those designed to mitigate such vulnerabilities. For example, integrating a diffusion model with an image classifier might address adversarial examples, with the model card specifying this capability as a system property rather than an attribute of the individual model (Carlini et al. 2022b). In this scenario, adversarial examples would be considered “in-scope” and reportable. Similarly, in the case of an LLM, deploying a hate speech detection model alongside the LLM could prevent the generation of hate speech, with the LLM model card noting the potential for hate speech, while the system card indicates that hate speech is effectively blocked. Another example includes Dalle2’s use of demographic descriptors to enhance output diversity (OpenAI 2022).
@@ 4.3 Automatic Verification
Model errors are anticipated to constitute a significant proportion of reports in any database of AI system issues. A structured reporting mechanism, such as a telemetry system, offers several advantages to vendors. Primarily, it facilitates seamless integration of reported issues into the model’s training and testing datasets. This approach also ensures that the reported behavior is verifiable and reproducible—a critical consideration given the potential for daily fluctuations due to model updates and continuous training processes.
The majority of substantive reports, as outlined in the background section, include demonstrative samples of the problematic behavior. Therefore, any system supporting Coordinated Flaw Disclosure should incorporate functionalities for automatic external verification and integration into the vendor’s internal testing dataset. This approach significantly reduces the vendor’s resource expenditure in issue remediation and provides a mechanism for ongoing verification of issue resolution.
Upon acceptance of a reported issue, the system should preserve the supporting data and implement periodic re-verification protocols. This is particularly crucial as issue resolution in one model iteration does not guarantee its absence in subsequent versions, potentially leading to regression and issue recurrence. By maintaining a robust verification system, vendors can ensure continued model integrity and swiftly address any re-emerging issues.
@@
4.4 Dynamic Scope Expansion for Common Use
For general-purpose models555also called foundation models or frontier models, some of these models are additionally open-weight with permissive licensing and have led to an ecosystem of downstream use cases, such as Meta’s LLAMA family of models (https://llama.meta.com/). with potentially limitless applications, we propose a “Common Use” clause in the model card. This clause allows for the dynamic expansion of the model’s intent and scope based on widespread adoption of certain use cases, even if they were not originally defined.
Initially, individual companies should strive to identify and incorporate common uses of their models. However, as the field matures, this process could evolve into a centralized effort, potentially managed by a cross-vendor coalition or an independent body. Such an approach could lead to the development of a Common Use Enumeration (CUE), complementing existing systems like the Common Weakness Enumeration (CWE).
To implement this approach:
  1. 1.
    Define a threshold for “common use”: Establish quantitative criteria (e.g., X% of users or Y number of deployments) to determine when a use case becomes ’common’.
  2. 2.
    Regular review process: Implement a periodic review (e.g., quarterly or bi-annually) to identify emerging common uses of the model.
  3. 3.
    Scope expansion: For uses meeting the defined threshold, expand the model card to include these as in-scope for flaw reporting and mitigation.
  4. 4.
    Allow reporting: Enable the submission of issues related to these expanded use cases through the CFD process.
The Common Use clause acknowledges that while general-purpose models may have vast potential applications, safety and responsible development require some boundaries. By allowing for dynamic scope expansion, we create a framework that can evolve with the actual use patterns of the model, enhancing the overall effectiveness of the CFD process. As the AI field continues to advance, the development of a centralized, independent CUE system could further streamline this process and promote industry-wide consistency in addressing common uses of AI models. We further discuss a CUE system in the Future Work section of this paper.
@@ 5.2 CFE Issuance Process
5.3 Adjudication Process for Edge Cases
While the CFD process primarily focuses on issues within the defined scope and intent of a model, the dynamic nature of AI systems, particularly general-purpose models, necessitates a mechanism for addressing edge cases. These edge cases often arise from common uses that were not initially anticipated in the model card. To address this, we propose an extended adjudication process:
  1. 1.
    Initial Triage: When a report is submitted that falls outside the stated intent but appears to be within common use, it is flagged for special review.
  2. 2.
    Common Use Assessment: The adjudication panel first determines if the reported use case meets the criteria for ’common use’ as defined in the Dynamic Scope Expansion clause (Section 4.4).
  3. 3.
    Harm Evaluation: If the use case is deemed common, the panel assesses the severity and prevalence of the reported harm.
  4. 4.
    Vendor Consultation: The panel consults with the vendor to understand any technical constraints or considerations related to the reported issue.
  5. 5.
    Decision: Based on these factors, the panel may:
    • Recommend immediate inclusion of the issue in the CFD process
    • Suggest updating the model card to include this use case in future versions
    • Advise on interim mitigation strategies
    • Determine that the issue, while outside of scope, warrants public disclosure due to its potential impact
  6. 6.
    Follow-up: The panel ensures that any decisions are implemented, including updates to the model card, CFD database, or public disclosures as necessary.
This extended process balances the need for clear boundaries with the reality of evolving AI applications, providing a structured approach to handling edge cases that may otherwise fall through the cracks of traditional disclosure processes.
@@ 6 Conclusion
This paper introduces a Coordinated Flaw Disclosure (CFD) framework for machine learning and AI systems, addressing algorithmic flaws through transparent intent and scope definitions via extended Model Cards. The CFD framework balances submitter and vendor interests, fostering trust and transparency through structured issue reporting governance. By considering both quantitative and qualitative evidence, it captures a holistic view of AI system flaws. We believe widespread adoption of this framework can bolster public trust in AI systems by promoting clear processes and collaborative engagement. As machine learning evolves, the CFD framework offers a structured, adaptable approach to navigate the changing landscape of algorithmic flaws, benefiting both vendors and the public.
6.1 Future Work
While our proposed CFD framework lays the groundwork for improving AI accountability and transparency, there are several areas where further development and real-world testing are needed. We outline two key areas for future work: an implementation pilot to test the CFD framework in practice, and the development of a Common Use Enumeration system to standardize the tracking of AI model applications.
Implementation Pilot at DEF CON 32
In collaboration with key stakeholders from government, industry, freelance security researchers, and policy experts, we are organizing the Generative Red Team 2 (GRT2) [^7] https://grt.aivillage.org/announcement event at DEF CON 32. This event serves as a comprehensive test of the CFD framework’s robustness and applicability in real-world scenarios. Key aspects of GRT2 that align with and test our CFD framework include:
  • Model Cards: Participants receive detailed model cards specifying the intent and scope of the target AI system, mirroring our proposed extended documentation approach.
  • Comprehensive Flaw Reports: Using the UK AI Safety Institute’s Inspect AI framework [^8] https://inspect.ai-safety-institute.org.uk/, participants prepare statistically valid reports on model flaws, moving beyond single-example vulnerabilities.
  • Dedicated Submission Platform: Reports are submitted through a platform, demonstrating the feasibility of specialized CFD infrastructure.
  • Vendor Interaction: AI system vendors review and respond to submitted reports, simulating real-world disclosure dynamics.
  • Independent Adjudication: A panel of independent experts mediates disputes between submitters and vendors, testing the efficacy of our proposed adjudication process.
  • Transparency: All data generated during the event will be made public, aligning with our goals for increased transparency in AI flaw disclosure.
GRT2 is designed not only to validate the CFD framework but also to stress-test it, identifying potential weaknesses and areas for improvement. The results and lessons learned from GRT2 will be invaluable in refining the CFD framework.
Common Use Enumeration (CUE) System
To address the challenge of tracking and standardizing common uses of AI models, we propose the development of a Common Use Enumeration (CUE) system. This system would serve as a centralized repository for documenting and categorizing the various applications of AI models, particularly focusing on general-purpose or foundation models that may have a wide range of unforeseen uses.
Proposed Structure of the CUE System
  1. 1.
    Hierarchical Classification: The CUE system would organize AI model uses into a hierarchical structure, similar to the Common Weakness Enumeration (CWE) system. This would allow for broad categories (e.g., “Natural Language Processing”, “Computer Vision”) with increasingly specific subcategories.
  2. 2.
    Unique Identifiers: Each documented use case would receive a unique identifier (e.g., CUE-001) for easy reference and tracking.
  3. 3.
    Metadata Fields: Each entry would include relevant metadata such as description of the use case, associated model types, potential risks or ethical considerations, known implementations or case studies, and related CUEs or CWEs.
  4. 4.
    Version Control: The system would maintain version history to track how use cases evolve over time.
Management and Governance
We propose the establishment of a multi-stakeholder governing body to oversee the CUE system, modeled after established multi-stakeholder processes in the CVE ecosystem. This body would manage a standardized submission process, regular updates, and ensure open access to the CUE database. Vendors could integrate CUE identifiers into their model cards, use the system for reporting new uses and risk assessment, and access an API for integrating CUE data into their development and monitoring tools.
The development of a Common Use Enumeration system represents a significant undertaking that would require collaboration across the AI industry, academia, and regulatory bodies. However, such a system could greatly enhance our ability to track, understand, and responsibly manage the diverse applications of AI models in society.
@@ 6.2 Limitations
While the CFD framework presents a significant step forward, several challenges remain:
  • Model Cards Standardization: The lack of standardized model cards poses a challenge for widespread CFD adoption. Establishing standards for model cards is crucial for effective implementation, as emphasized by the NIST AI Risk Management Framework (Singh 2021).
  • Dynamic ML Datasets: Tracking and managing evolving datasets used in model training presents complex challenges. Unlike software dependencies tracked by SBOMs (CISA 2021; Corman 2022), monitoring dynamic ML datasets lacks established tools. Recent incidents, such as the LAION-5B dataset retraction (Schuhmann et al. 2022; LAION.ai 2023; Thiel 2023), highlight the need for robust dataset management in machine learning.
  • Contextual Challenges: The context-specific nature of AI systems complicates the development of standardized weakness and use enumerations. For instance, while a system might be theoretically vulnerable to “Prompt Injection” (MITRE 2024b), its specific application might mitigate this risk. Identical ML models used for different purposes add complexity, making model documentation crucial for understanding relevant weaknesses and uses across various contexts.
Addressing these limitations requires ongoing collaboration between AI developers, security researchers, and domain experts to create flexible, adaptive systems that accurately capture the nuances of AI weaknesses and uses across diverse applications.